# Electric dipole polarizability constraints on neutron skin and symmetry energy

### v1

@@ 6.2 Constraints based on ab initio models
Figure 11: (A) Experimental DP in 48Ca (blue band) and predictions from ab initio results based on $\chi$EFT interactions (green triangles) and DFT calculations (red squares). The green and black bars indicate the ab initio prediction selected to reproduce the 48Ca charge radius and the range of DP predictions from Ref. [101] simultaneously consistent with the DP in 68Ni, 120Sn and 208Pb, cf. Fig. 11(D). (B) Correlation of the experimental DP (green band) and the charge radius (black band) in 68Ni with comparison to ab initio coupled-cluster calculations up to 2p-2h (dashed crosses) and 3p-3h excitations (full crosses). The dashed and full lines and corresponding error bands result from linear fits to the theoretical resuts. (C) Correlation of the experimental DP in 40Ca and 48Ca in comparison with ab initio coupled-cluster calculations including 3p-3h excitations (crosses and purple uncertainty band). Figures taken from (A) Ref. [16], (B) [50] and (C) [31], where the original references can be found.
The EDF results tend to be somewhat high compared to experiment. The ab initio results show a significant dependence on the chosen interaction, but it can be well approximated by a linear dependence. This allows in principle to derive boundaries on the neutron skin thickness and the symmetry energy. However, while the ab initio results shown were truncated in the coupled-cluster expansion at the 2p-2h level, subsequent work [77] demonstrated that inclusion of 3p-3h correlations lowers the $\alpha_{\rm D}$ values by $10-20$%. The refined results in 48Ca are plotted in Fig. 11(C) against corresponding calculations for 40Ca [31]. A high correlation similar to the DFT results shown in Fig. 10(A) is observed. The pink uncertainty band from the ab initio results overlaps with the crossing of the experimental $1\sigma$ error bands. In particular, the NNLOsat interaction interaction [27] accurately describing binding energies and radii of nuclei up to 40Ca as well as the saturation point of symmetric nuclear matter now reproduces both DP values. The importance of including 3p-3h correlations has also been demonstrated in a recent measurement of the 68Ni charge radius [50]. Figure 11(B) illustrates the improvement in describing the correlation between the charge radius and $\alpha_{\rm D}$ [103] when going from the 2p-2h level (light blue band) to the inclusion of 3p-3h correlations (dark blue band).
@@ 6.3 Tension between polarizability and parity-violating elastic electron scattering in 208Pb
While in 48Ca there is fair agreement between the neutron skin thickness and symmetry energy properties derived from the different experiments, the parity-violating elastic electron scattering experiment on 208Pb [1] finds a much larger neutron skin $r_{\rm skin}=0.28(7)$ fm than most other work. Accordingly, an extraction of symmetry energy parameters based on the correlations established in DFT (see Sec. 4.1) leads to large symmetry energy values $J=38(5)$ MeV and $L=106(37)$ MeV in contradiction to limits derived from astrophysical observations of neutron star radii and masses as well as the tidal deformability of neutron star mergers [62]. All astrophysical constraints point towards a softer EOS. This has led to speculations about a phase transition at intermediate densities [40].
@@
Because of the strong correlation between $r_{\rm skin}$ and $\alpha_{\rm D}$ for a given nucleus and $r_{\rm skin}$ values of different nuclei in DFT models, Reinhard et al. [96, 97] investigated whether it is possible to construct a DFT interaction capable of simultaneously describing the data for 48Ca and 208Pb. The analysis was based on representative families of non-relativistic and relativistic functionals. The isovector properties of EDFs are typically not well constrained by the input data used to fit the model parameters. As illustrated in Fig. 12(A) for the case of 208208{208}208Pb, it is possible to vary the symmetry energy parameters – and thereby the predicted $r_{\rm skin}$ and $\alpha_{\rm D}$ over a fairly large range maintaining comparable description of ground-state properties [96]. Figure 12(B) [97] demonstrates that the polarizabilities and the neutron skin thickness of 48Ca could be consistently described, but it was impossible to construct an EDF simultaneously accounting for the neutron skin thickness of 208Pb extracted from the PREX experiment [2]. Similar conclusions were drawn in Refs. [90, 134]. In another recent attempt [94], a DFT interaction reasonably accounting for the measured parity-violating asymmetries in both the PREX and CREX experiments was constructed, but at the expense of unusual properties of the symmetry energy curvature and a very strong isovector coupling leading to density fluctuations in the nuclear interior.
@@ 6.4 Volume and surface contributions to the symmetry energy
Another way of extracting properties of the symmetry energy is a study of the mass dependence of the DP. A simple power law $\sigma_{-2}\propto A^{5/3}$ based on a model of two interpenetrating fluids has been given by Migdal, where $\sigma_{2}$ denotes the second inverse moment of the photoabsorption cross sections and $\sigma_{2}\simeq\alpha_{\rm D}$ in units of mb/MeV (Ref. [86] and Refs. therein). A proportionality constant $2.4\times 10^{-3}$ has been determined by Orce [85] from a fit to $(\gamma,xn)$ data over a wide mass range. Figure 13 [125] shows a comparison with a combined data set of $\alpha_{\rm D}$ measurements in light nuclei [5] with then available (2016) data from relativistic Coulomb excitation for heavier nuclei as green short-dashed line. (Note that results for $A<12$ from Ref. [5] are neglected because the hydrodynamical picture is highly questionable and corrections due to the magnetic polarizability are large [55] for these very light nuclei.) Results are severely underestimated in lighter nuclei where charged-particle decay dominates. The mass dependence is reasonably described for larger masses but the proportionality coefficient of Ref. [85] is too low, since additional contributions from the strength below neutron threshold, as discussed in Sec. 5.3, need to be considered.
Figure 13: Experimental DP for a set of nuclei as a function of mass number (full squares). The green and blue lines are fits with original Migdal model (Eqs. (1) and (2) in Ref. [85]). The black lines are fits allowing for a surface term of the symmetry energy including (dashed-dotted) and excluding (full) the data point for 12C. The red line shows a fit with the prediction of Ref. [109] using the ”$\mu_{n}$” approach. Figure taken from Ref. [125], where the original references can be found.
@@
Here $\kappa=S_{s}/S_{v}$, and $S_{s}$ and $S_{v}$ denote the surface and volume coefficients of the symmetry energy, respectively. The numerical coefficient is obtained from Migdal’s approach. A fit with $S_{s},S_{v}$ parameters from binding energies of isobaric nuclei [116] shown in Fig. 13 as blue long-dashed line still underestimates the lower-mass data. Parameters of the study of Ref. [109] provide a better description (dotted red line). Results of a free fit of Eq. 5 crucially depend on the inclusion (dotted-dashed black line) or exclusion (solid black line) of the 12C data point. The latter provides a better fit with $S_{v}=25.6(8)$ MeV, $\kappa=1.66(5)$ [125] close to Ref. [109]. $S_{v}$ can be interpreted as $J$, but measured at about $2/3$ of the saturation density [22, 99].
@@ 7 Conclusions and outlook
While the mass depedence of the DP is reasonable covered by the available data, future work should explore other degrees of freedom like the variation of neutron excess along isotopic chains and the role of deformation. Furthermore, the experimental uncertainties of the DP for key nuclei can be improved by the availability of independent measurements as illustrated in Fig. 6. New high-brilliance LCBS photon beam facilities are under construction at the Extreme Light Infrastructure - Nuclear Physics (ELI-NP) in Bucarest [33, 114] and the Shanghai Laser Electron Gamma Source (SLEGS) at the Shanghai Synchrotron Radiation Facility [130]. Combined with advanced techniques for neutron detection [35], these facilities promise a new quality of precision for $(\gamma,xn$) experiments.

### v3

@@ 6.2 Constraints based on ab initio models
Figure 11: (A) Experimental DP in 48Ca (blue band) and predictions from ab initio results based on $\chi$EFT interactions (green triangles) and DFT calculations (red squares). The green and black bars indicate the ab initio prediction selected to reproduce the 48Ca charge radius and the range of DP predictions from Ref. [102] simultaneously consistent with the DP in 68Ni, 120Sn and 208Pb, cf. Fig. 11(D). (B) Correlation of the experimental DP (green band) and the charge radius (black band) in 68Ni with comparison to ab initio coupled-cluster calculations up to 2p-2h (dashed crosses) and 3p-3h excitations (full crosses). The dashed and full lines and corresponding error bands result from linear fits to the theoretical resuts. (C) Correlation of the experimental DP in 40Ca and 48Ca in comparison with ab initio coupled-cluster calculations including 3p-3h excitations (crosses and purple uncertainty band). Figures taken from (A) Ref. [17] and (B) [51], where the original references can be found. Figure 11(C) is taken from Ref. [32] but modified to include an estimate of the theoretical uncertainties shown as purple band.
The EDF results tend to be somewhat high compared to experiment. The ab initio results show a significant dependence on the chosen interaction, but it can be well approximated by a linear dependence. This allows in principle to derive boundaries on the neutron skin thickness and the symmetry energy. However, while the ab initio results shown were truncated in the coupled-cluster expansion at the 2p-2h level, subsequent work [78] demonstrated that inclusion of 3p-3h correlations lowers the $\alpha_{\rm D}$ values by $10-20$%. The refined results in 48Ca are plotted in Fig. 11(C) against corresponding calculations for 40Ca [32]. A high correlation similar to the DFT results shown in Fig. 10(A) is observed. The purple uncertainty band from the ab initio results overlaps with the crossing of the experimental $1\sigma$ error bands. In particular, the NNLOsat interaction interaction [28] accurately describing binding energies and radii of nuclei up to 40Ca as well as the saturation point of symmetric nuclear matter now reproduces both DP values. The importance of including 3p-3h correlations has also been demonstrated in a recent measurement of the 68Ni charge radius [51]. Figure 11(B) illustrates the improvement in describing the correlation between the charge radius and $\alpha_{\rm D}$ [104] when going from the 2p-2h level (light blue band) to the inclusion of 3p-3h correlations (dark blue band).
@@ 6.3 Tension between polarizability and parity-violating elastic electron scattering in 208Pb
While in 48Ca there is fair agreement between the neutron skin thickness and symmetry energy properties derived from the different experiments, the parity-violating elastic electron scattering experiment on 208Pb [1] finds a much larger neutron skin $r_{skin}=0.28(7)$ fm than most other work. Accordingly, an extraction of symmetry energy parameters based on the correlations established in DFT (see Sec. 4.1) leads to large symmetry energy values $J=38(5)$ MeV and $L=106(37)$ MeV in contradiction to limits derived from astrophysical observations of neutron star radii and masses as well as the tidal deformability of neutron star mergers [63]. All astrophysical constraints point towards a softer EOS. This has led to speculations about a phase transition at intermediate densities [41].
@@
Because of the strong correlation between $r_{\rm skin}$ and $\alpha_{\rm D}$ for a given nucleus and $r_{\rm skin}$ values of different nuclei in DFT models, Reinhard et al. [97, 98] investigated whether it is possible to construct a DFT interaction capable of simultaneously describing the data for 48Ca and 208Pb. The analysis was based on representative families of non-relativistic and relativistic functionals. The isovector properties of EDFs are typically not well constrained by the input data used to fit the model parameters. As illustrated in Fig. 12(A) for the case of 208Pb, it is possible to vary the symmetry energy parameters – and thereby the predicted $r_{skin}$ and $\alpha_{D}$ – over a fairly large range maintaining comparable description of ground-state properties [97]. Figure 12(B) [98] demonstrates that the polarizabilities and the neutron skin thickness of 48Ca could be consistently described, but it was impossible to construct an EDF simultaneously accounting for the neutron skin thickness of 208Pb extracted from the PREX experiment [2]. Similar conclusions were drawn in Refs. [91, 136]. In another recent attempt [95], a DFT interaction reasonably accounting for the measured parity-violating asymmetries in both the PREX and CREX experiments was constructed, but at the expense of unusual properties of the symmetry energy curvature and a very strong isovector coupling leading to density fluctuations in the nuclear interior.
@@ 6.4 Volume and surface contributions to the symmetry energy
Another way of extracting properties of the symmetry energy is a study of the mass dependence of the DP. A simple power law $\sigma_{-2}\propto A^{5/3}$ based on a model of two interpenetrating fluids has been given by Migdal, where $\sigma_{-2}$ denotes the second inverse moment of the photoabsorption cross sections and $\sigma_{2}\simeq\alpha_{\rm D}$ in units of mb/MeV (Ref. [87] and Refs. therein). A proportionality constant $2.4\times 10^{-3}$ has been determined by Orce [86] from a fit to $(\gamma,xn)$ data over a wide mass range. Figure 13 [127] shows a comparison with a combined data set of $\alpha_{D}$ measurements in light nuclei [5] with then available (2016) data from relativistic Coulomb excitation for heavier nuclei as green short-dashed line. Note that results for $A<12$ from Ref. [5] are neglected because the hydrodynamical picture is highly questionable and corrections due to the magnetic polarizability are large [56] for these very light nuclei. Results are severely underestimated in lighter nuclei where charged-particle decay dominates. The mass dependence is reasonably described for larger masses but the proportionality coefficient of Ref. [86] is too low, since additional contributions from the strength below neutron threshold, as discussed in Sec. 5.3, need to be considered.
Figure 13: Experimental DP for a set of nuclei as a function of mass number (full squares). The green and blue lines are fits with original Migdal model (Eqs. (1) and (2) in Ref. [86]). The black lines are fits allowing for a surface term of the symmetry energy including (dashed-dotted) and excluding (full) the data point for 12C. The red line shows a fit with the prediction of Ref. [110] using the "$\mu_{n}$" approach. Figure taken from Ref. [127], where the original references can be found.
@@
Here $\kappa=S_{s}/S_{v}$, and $S_{s}$ and $S_{v}$ denote the surface and volume coefficients of the symmetry energy, respectively. The numerical coefficient is obtained from Migdal’s approach. A fit with $S_{s},S_{v}$ parameters from binding energies of isobaric nuclei [118] shown in Fig. 13 as blue long-dashed line still underestimates the lower-mass data. Parameters of the study of Ref. [110] provide a better description (dotted red line). Results of a free fit of Eq. (5) crucially depend on the inclusion (dotted-dashed black line) or exclusion (solid black line) of the 12C data point. The latter provides a better fit with $S_{v}=25.6(8)$ MeV, $\kappa=1.66(5)$ [127] close to Ref. [110]. $S_{v}$ can be interpreted as $J$, but measured at about $2/3$ of the saturation density [23, 100].
@@ 7 Conclusions and outlook
While the mass dependence of the DP is reasonable covered by the available data, future work should explore other degrees of freedom like the variation of neutron excess along isotopic chains and the role of deformation. Furthermore, the experimental uncertainties of the DP for key nuclei can be improved by the availability of independent measurements as illustrated in Fig. 6. New high-brilliance LCBS photon beam facilities are under construction at the Extreme Light Infrastructure - Nuclear Physics (ELI-NP) in Bucarest [34, 115] and the Shanghai Laser Electron Gamma Source (SLEGS) at the Shanghai Synchrotron Radiation Facility [132]. Combined with advanced techniques for neutron detection [36], these facilities promise a new quality of precision for $(\gamma,xn$) experiments.